Cost of Equity

Radon Homes's current EPS is $6.57. It was $4.38 5 years ago. The company pays out 60% of its earnings as dividends, and the stock sells for $30.

  1. Calculate the historical growth rate in earnings. (Hint: This is a 5-year growth period.) Round your answer to two decimal places.
         %
  2. Calculate the next expected dividend per share, D1 (Hint: D0 = 0.60($6.57) = $3.94). Assume that the past growth rate will continue. Round your answer to the nearest cent.
    $   
  3. What is Radon's cost of equity, rs? Round your answer to two decimal places.
         %
Answer #1


solutgon: FV=PV*(1+g)^n 6.57=4.38*(1+g)^5 (1+g)^5=6.57/4.38 (1+g)^5=1.5 1+g=1.5^(1/5) 1+g=1.084472 g=1.084472-1 FV Future val